### v2

# Date of publication: 24 April 2025

Abstract
Daily Activity Recordings for artificial intelligence (DARai, pronounced /Dahr-ree/), is a multimodal, hierarchically annotated dataset constructed to understand human activities in real-world settings. DARai consists of continuous scripted and unscripted recordings of 50 participants in 10 different environments, totaling over 200 hours of data from 20 sensors including multiple camera views, depth and radar sensors, wearable inertial measurement units (IMUs), electromyography (EMG), insole pressure sensors, biomonitor sensors, and gaze tracker. To capture the complexity in human activities, DARai is annotated at three levels of hierarchy: (i) high-level activities (L1) that are independent tasks, (ii) lower-level actions (L2) that are patterns shared between activities, and (iii) fine-grained procedures (L3) that detail the exact execution steps for actions. The dataset annotations and recordings are designed so that $22.7\%$ of L2 actions are shared between L1 activities and $14.2\%$ of L3 procedures are shared between L2 actions. The overlap and unscripted nature of DARai allows counterfactual activities in the dataset. Experiments with various machine learning models showcase the value of DARai in uncovering important challenges in human-centered applications. Specifically, we conduct unimodal and multimodal sensor fusion experiments for recognition, temporal localization, and future action anticipation across all hierarchical annotation levels. To showcase the shortcomings of individual sensors, we conduct domain-variant experiments that are possible because of DARai’s multi-sensor and counterfactual activity design setup. The code, documentation, and dataset is available at the dedicated DARai website.
Keywords: Multimodal Fusion, Temporal Sequence Modeling, Cross-View Domain Adaptation, Multi-sensor Integration, Hierarchical Learning, Hierarchical Activity Recognition, Time-Series Analysis, Real-World Environments, Action Anticipation, Action Segmentation
1 Introduction
Figure 1: DARai multimodal data collected to capture and reflect counterfactual variations in activities across different modalities. Hierarchical structure of DARai data, demonstrated using an example activity: Moving an Object. Two variations are shown, moving a heavy box (blue) and a light box (green), along with their corresponding lower hierarchy levels. For instance, variations in insole pressure and forearm muscle interactions are observed when lifting or carrying a box.
Challenges in visual data understanding
Visual information has traditionally driven research in understanding human activities. Progress in computer vision techniques for tasks such as recognition and scene understanding (Kaur et al., 2024), image classification and detection (Carion et al., 2020), image segmentation (Quesada et al., 2024), video action recognition (Ma et al., 2019), and traffic surveillance (Kokilepersaud et al., 2023b) are centered around visual data. However, in real-world settings, visual processing methods face significant practical and ethical constraints, including partial occlusions (Gunasekaran and Jaiman, 2023), acquisition and environmental errors (Temel et al., 2017b, a, 2019; Prabhushankar and AlRegib, 2022), sensitivity to angles and viewpoints (Ozturk et al., 2024; Dong et al., 2022; Garg et al., 2022), privacy concerns (Mujirishvili et al., 2023; Ravi et al., 2024), and the impracticality of maintaining fixed or unobtrusive camera perspectives. In contrast, widely available wearable sensors like inertial measurement units (IMUs), electromyography (EMG), and pressure insoles capture body-centric signals unaffected by camera placement. Furthermore, they provide additional information than visual data alone.
Value in multimodal and hierarchical data acquisition
Consider the activity of Moving an object demonstrated in Figure 1. This simple activity can differ substantially in muscle interaction, foot pressure distribution, and physiological responses depending on whether the object is light or heavy. The differences in moving a heavy or light object may be visually imperceptible but are clearly evident in insole pressure data as demonstrated in Figure 1. Therefore, integrating various data modalities help mitigate the limitations of a single modality approach (Lin et al., 2023). Such multimodal integration becomes essential given that human activities follow a hierarchical structure. The activity hierarchy starts with a specific goal which we term as Level 1 or L1 Activity. Moving an object in Figure 1 is an L1 activity. Next, intermediate actions, termed Level 2 or L2 are planned. These actions include pick up, carry and put down in Figure 1. Finally, the specific procedures, including bend and lift, that complete each action are executed. Human-centric datasets require capturing multimodal and nested L1, L2, and L3 hierarchies that reflect both high-level and fine-grained level sensor signals and temporal organization (Robinson and Sloutsky, 2010).
Dataset requirements for daily activity understanding
Research on human-centered applications like robotics (Chen et al., 2021), assistive technologies (Bouchabou et al., 2021), health monitoring (Wang et al., 2023), and disease diagnosis (Prabhushankar et al., 2022) utilize visual data and additional sparse data modalities. These additional modalities include skeleton data (Wang and Yan, 2023), clinical and biomarker data (Kokilepersaud et al., 2023a), and more recently language models (Damen et al., 2018) among others. However, commonly used datasets in daily activity and action understanding emphasize single-modality (Byrne et al., 2023) or scripted scenarios (Stein and McKenna, 2013) and rarely incorporate hierarchical annotations across multiple sensors (Liu et al., 2019). Addressing these gaps by designing datasets that feature continuous, unscripted recordings from both visual and non-visual sensors can enable researchers to tackle significant challenges related to privacy, multimodal learning strategies, and hierarchical activity recognition, thus promoting robust, real-world human activity understanding.
In this paper, we propose the Daily Activity Recordings for artificial intelligence (DARai) dataset. DARai is an open source, multimodal, hierarchical, and continuously recorded dataset. DARai marks the largest available dataset interms of the number of sensors and data modalities. The contributions of DARai are listed as follows.
  1. 1.
    DARai is a comprehensive daily activity dataset consisting of $200$ hours of recordings across $20$ data modalities from $12$ sensors. DARai is collected from $50$ participants in five different kitchens and living spaces, totaling 10 unique environments.
  2. 2.
    DARai is annotated to have $160$ classes across three levels of hierarchies - activities (L1), actions (L2), and procedures (L3). Furthermore, an action level language description of the L1 data is provided.
  3. 3.
    The hierarchical structure and continuously recorded data allows construction of counterfactual action scenarios and long-range temporal dependencies, thereby showcasing the utility of specific sensors under different conditions.
  4. 4.
    DARai is anonymized, preprocessed and curated for machine learning applications and is open sourced at IEEE Dataport (Kaviani et al., 2024).
  5. 5.
    Benchmarks for activity, action, and procedure recognition, action and procedure temporal segmentation, and action anticipation are provided on DARai and the codes are available on the project page.
Challenges in DARai data collection
Gathering continuous, real-world data across multiple environments using various sensor platforms is a significant challenge. Notably, each of the $12$ sensors used in DARai and listed in Table 3, have different sampling rates and timestamps’ protocol, making device synchronization challenging. This is addressed by implementing a multi-threaded system with a global timestamp signal to align all start and end time of data streams effectively. More information on sensor setup and calibration is provided in Appendix  LABEL:Appendix:_Sensor-Setup. We achieve a synchronization drift of 1 ms in 24-hour timespan. Depth sensors overheat during long recording sessions, reducing depth estimation quality and requiring periodic cool downs. Wearable sensors are connected wirelessly, and sometimes encounter connectivity drops. Such connectivity issues are overcome with post-recording recovery from device internal memory, thereby necessitating using sensors that provide significant internal memory. A further challenge is the large volume of data. Each participant in a session (spanning between 30 minutes to 1 hour) generates up to 300 GB of data. This necessitates local storage before transferring processed subsets to the servers. Details on data size and format are available in Appendix  LABEL:Appendix:_Structure.
The novelty of DARai is compared against existing datasets in Section 2. The specifics of data collection, annotation, and processing are described in Section 3, followed by machine learning experiment setup and benchmarks in Sections 456, and 7.
2 Related Work
In this section, we examine $28$ multimodal and hierarchical datasets to establish the novelty of DARai dataset.
Multimodal Datasets
Table 1: Datasets Comparison Summary
Setup Dataset # Classes # Subjects # Modality Activities Multiple
Environment
Scripted Activities CMU-MMAC (De la Torre et al., 2009) 5 18 4 Food preparation No
UT-Kinect (Xia et al., 2012) 10 10 3 Daily arm motions No
Multiview 3D Event (Wang et al., 2014) 8 8 3 Daily actions No
UTD-MHAD (Chen et al., 2015) 27 8 3 Exercise, Arm motions No
Stanford ECM (Nakamura et al., 2017) 24 10 3 Physical activities No
Daily Intention (Wu et al., 2017) 34 12 2 Daily actions No
UI-PRMD (Vakanski et al., 2018) 10 10 3 Physical rehabilitation No
NTU RGB+D 120 (Liu et al., 2019) 120 106 4 Indoor activities No
MMAct (Kong et al., 2019) 37 20 3 Daily life Yes
UESTC RGB-D (Ji et al., 2019) 118 40 2 Aerobic exercise No
OREBA (Rouast et al., 2020) 4 102 3 Eating No
LaRa (Niemann et al., 2020) 8 14 3 Logistic and packaging No
BON (Tadesse et al., 2021) 18 25 1 Office activities Yes
ActionSense (DelPreto et al., 2022) 20 10 8 Kitchen activities No
StrokeRehab (Kaku et al., 2022) 24 41 2 Stroke rehabilitation No
MPHOI-72 (Qiao et al., 2022) 13 5 2 Human and object Interaction No
Hybrid Breakfast Actions (Kuehne et al., 2014b) 10 52 1 Food preparation No
DARai (2024) 160 50 20 Daily activity, Office tasks, Kitchen and Household activities Yes
Unscripted Activities Epic kitchen-100 (Damen et al., 2018) 97 37 3 Kitchen activities No
Ego4D (Grauman et al., 2022) Varies* 931 4 Daily activities
(household, outdoor, workplace, etc.)
Yes
Ego-Exo4D (Grauman et al., 2023) Varies* 839 6 Skilled human activities
(sports, music, dance, etc.)
Yes
CAP (Byrne et al., 2023) 512 780 1 Indoor activities Yes
*Varies indicates that the exact set of activities present in the dataset is unknown and can only be determined by accessing the annotations.
Several widely used multimodal datasets are shown in Table 1, organized by their data collection setups: scripted activities, unscripted activities, or a hybrid approach. The data collection setup influences trade-offs between the number of activity classes, subjects, sensors setup and environments, as well as activity categories.
  • Scripted Activities: Participants receive instructions to perform scripted activities. Generally, participants carry out similar activities. Such setup limits the variety in activity scenarios and environments, enabling the use of stationary and cumbersome sensors (e.g., motion capture suits).
  • Unscripted Activities: Participants act freely without a scripted set of activities. Such a set-up results in unscripted or "in-the-wild" data. This setup often results in diverse set of activities and environments. Though it is difficult to maintain both consistency and variety in the sensor setup under these conditions.
  • Hybrid: Participants are instructed to perform all or a subset of high-level activities from a predefined list while acting freely without a scripted plan for the lower level intermediate steps. Hybrid setups allow for more natural recordings of human activity in diverse environments while ensuring a diverse set of activities and supporting larger sensor setup.
In addition to the categorization provided in the columns of Table 1, DARai includes activities of varying lengths, from 30 seconds to over 5 minutes. This is opposed to existing datasets that provide short activity clips, without accounting for long-term temporal dependencies.
Hierarchical Datasets
Table 2: Comparison of Taxonomy in Hierarchical Datasets
Setup Dataset Hierarchy Structure Hierarchy Construction Method
Post-Collection Hierarchy Design 50 Salads (Stein and McKenna, 2013) High-level activity, Low-level activity Annotator decision
Breakfast Actions (Kuehne et al., 2014b) Coarse actions, Fine-grained actions Sentences from audio transcription
YouCook2 (Zhou et al., 2018b) Activity , Steps Sentences from subtitles
FineGym (Shao et al., 2020) Event, Set, Element Annotator decision, Decision-tree based process
MoMa (Luo et al., 2021) Activity , Sub-Activity, Atomic Actions Annotator decision
FineSports (Xu et al., 2024) Action, Sub-Action Annotator decision, MixSort-OC ((Cui et al., 2023))
Ego4D Goal-Step (Song et al., 2024) Goal, Steps, Sub-Steps Annotator decision, wikiHow
Predefined Hierarchy Design Assembly 101 (Sener et al., 2022) Coarse actions, Fine-grained actions Verb and noun from activity labels
DARai (2024) Activity, Action, Procedure Predefine decomposition rules
Unlike object detection and image classification datasets, which benefit from standard taxonomies like WordNet (Miller, 1995), human activity datasets often lack a widely adopted hierarchical or semantic framework. Existing datasets group activities under broad scenarios (e.g., daily tasks, exercise, kitchen activities) without defining multi-level hierarchies beyond activity categories. Table 2 surveys several datasets that offer explicit hierarchies and categorized by hierarchy setup. Post-collection hierarchy design implies that hierarchies are constructed at the annotation level after data is collected. In contrast, predefined hierarchy design allows data collection with hierarchies, inspired by activity planning and reasoning by humans. Predefined hierarchies allow variations in certain activities through multiple instances of the same activity with minor changes. For example, watching different TV programs, moving objects of varying sizes and weights, using a specific recipe to prepare food, or preparing food without a recipe are all counterfactual activities. DARai consists of these counterfactual activities performed by the same participant. Predefined hierarchical structure supports more detailed analyses of complex tasks by breaking them down into smaller actions and procedures (Sloutsky, 2010).
3 Dataset Description and Annotation
3.1 Data collection process and considerations
Table 3: Sensors and Modalities Information
Category Sensor Data Modalities Views
3rd POV Kinect Cameraazu (2023) RGB, Depth, IR 2
Lidar CameraCorporation (2021) RGB, Depth, Depth Confidence, IR 2/1
1st POV Microphone Audio 1
Wearable IMU Wrist acceleration, gyro and magnetic Field 2
Wearable EMG Forearm Muscle Signal 2
Eye Tracker RGB, Audio, Head IMU and Gaze 1
Insole Pressure Total foot pressure, Multi zone pressure 1
Wearable Bio monitors Respiration rate, ECG, Heartbeat rate and intervals 1
Ambient and Remote Stationary IMU Surface vibration 2
Radar Doppler cubes 1
Environmental Sensors CO2, Humidity, Temperature, Light 1
BMI, Exhaustion Level Self Report Form Text N/A
Fifty participants (male and female) are recruited through a volunteer call and their activities are recorded in ten environments across five locations, including living spaces, home offices, and kitchens, with variations in lighting, time of day, air conditioning, and background noise. Each session includes counterfactual activity instances, where participants perform tasks with enforced variations, such as moving a heavy box instead of a light one. Some of these variations are detected by specific sensors like insole pressure sensors in Figure 1. All participants provide informed consent, are fairly compensated, and could opt out at any time, following an IRB-approved protocol that ensures participants’ privacy by removing personally identifiable information except for explicitly approved demonstration cases. Data is collected from 12 sensor devices, including Kinect, LiDAR, wearable IMUs on both hands, and stationary IMUs positioned on desks and floors in two environments, covering motion, physiological responses, and environmental conditions. Table 3 lists the sensors, their output modalities, and number of sensors used. Cameras are placed at different angles to capture standing and sitting postures, allowing natural movement rather than requiring participants to face a fixed camera, unlike other datasets (DelPreto et al., 2022; Liu et al., 2019; Qiao et al., 2022; Sener et al., 2022). After the recordings, participants complete two questionnaires: the first gathers information about their familiarity with the tasks, exhaustion levels, and BMI, while the second assesses their experience with the sensors. The survey results are provided in Appendix LABEL:Appendix:_Surveys, while additional technical details on sensor specifications, calibration, and sampling rates are in Appendix LABEL:Hardware, with a summary of sensor data modalities in Table LABEL:tab:sensors_spec.
3.2 Annotations
The annotation description of DARai hierarchy is provided below:
  • Level 1 (L1): Activities
    High-level, independent tasks that broadly describe actions without specifying detailed goals or variations. For instance, Moving an object without defining its size or weight. Figure 3 shows the hierarchy for the L1 activity Sleeping.
  • Level 2 (L2): Actions
    Actions are shorter segments that frequently occur across multiple activities. A complete instance of an activity consists of a sequence of these actions. The order of actions or inclusion of specific steps may vary. For example, the action Prep ingredients might be skipped if ingredients are already prepared. Actions alone may lack sufficient context to definitively identify the activity. For instance, the L2 action Prepare for Activity in Figure 3 could indicate either preparing to sleep or organizing the bed after waking, depending on temporal context.
  • Level 3 (L3): Procedures
    Procedures specify the detailed fine-grained context in which actions are performed, clarifying differences between individual instances of the same action. For instance, the L3 procedures for the action Prepare for Activity in Figure 3 include Lay down and Use blanket.
  • Natural language descriptions:
    Natural-language description provided by human annotators, describing activities and interactions and the scene explicitly at the L2 level.
Figure 2: A sequence of frames depicting an example of an annotated activity sample: L1 activity ’Sleeping’ with a sequence of L2 actions: ’Prepare for Activity’ and ’Rest’, and L3 procedures: ’Lay Down’ and ’Use Blanket’. Boundaries indicate the start and end times of each annotation level within the corresponding upper-level video sample.
Hierarchy Level Number of Unique Labels Example
$L\_1$: Activities 18 "Moving an Object"
$L\_2$: Actions 44 "Pick up object" , "Put down object"
$L\_3$: Procedures 98 "Bend" , "Lift" , "Put"
Figure 3: Number of unique labels at each level of hierarchy
Some activities in the dataset do not span over all levels of the hierarchy. In Figure 3, the L2 action Rest does not contain any L3 procedures. This flexibility allows both procedural activities such as Exercise or Playing a game, and non-procedural activities including Sleeping or Watching, to fit within the same hierarchy structure and definition. In contrast, other hierarchical datasets, including (Song et al., 2024; Sener et al., 2022; Zhou et al., 2018a; Bansal et al., 2022; Kuehne et al., 2014a), have only considered procedural activities in their design. We provide the number of unique labels at each level of the hierarchy, along with examples, in Table 3. Further details about the annotation workflow are provided in Appendix LABEL:Appendix:_Annotation.
3.3 Data Preprocessing
The DARai multisensory setup records over 200 hours of data across 20 modalities from 12 sensor devices, resulting in more than 20 TB of raw recordings. Appendix LABEL:Appendix:_Structure and Table LABEL:tab:sensors_spec in Appendix LABEL:Hardware detail the specific data formats, sampling rates, and resolutions for each sensor. The raw video data is anonymized by blurring participant faces. A standard imputation method is applied for missing values on hand EMG signals, insole pressure, and biomonitoring signals along with subject-specific normalization (Gasparini et al., 2022) for each data sample. The processed data are segmented into activity-level files based on L1 annotations, making the activity samples ready for machine learning and deep learning applications. The dataset is open-sourced and available on IEEE Dataport.
4 Experiment Setup for ML Applications
Figure 4: Overview of machine learning applications using the multimodal hierarchical DARai dataset. This figure highlights key aspects explored in this work, including temporal modeling, multimodal modeling, and hierarchical modeling, as well as potential applications for the machine learning community as direct or combined downstream tasks of these approaches.
As shown in Figure 4, DARai allows multimodal, hierarchical, and temporal modeling of human daily activities. In this paper, we choose three exemplar applications and demonstrate the challenges and utility of learning from human daily activity data in the following three sections. Section 5 evaluates the robustness of visual and wearable modalities under cross-view and cross-body settings, comparing fine-tuned and from-scratch deep learning models. Section 6 investigates unimodal and multimodal performance across the hierarchical levels of DARai for fine-grained activity recognition. Finally, Section 7 explores temporal dependencies in unscripted human activities, addressing both lower-level action localization and future activity anticipation. These tasks demonstrate DARai’s diverse downstream applications in machine learning.
Modality Dimension Resolution Information Provided
RGB, D 3 Channel, 1 Channel 15 fps, 30 fps Visual appearance, depth perception, spatial information
IMU 9 Channel 12 Hz, 200 Hz Tracking motion dynamics and hand interactions
EMG 1 Channel 2000 Hz, 4000 Hz Hand muscle activity , griping and holding
Insole 8 Channel 500 Hz Insole pressure distribution and weight balancing, moving patterns
Bio 5 Channel 1000 Hz Physiological signals
Gaze 2 Channel 200 Hz Visual focus tracking
Figure 5: Illustration of the unique information provided by various data modalities, highlighting the distinct insights each contributes to modeling human activity and state patterns.
Experiment Setup
In our experiments, we use a subset of the data modalities available in DARai, including RGB and Depth data from different camera views, IMU from both hands, EMG of both forearm muscles, insole pressure from feet, five biomonitoring signals, and gaze tracking data. Figure 5 summarizes the insights offered by each data modality, along with their dimensionality and resolution. The insights include motion dynamics, hand orientation, muscle activation patterns, foot pressure distribution and balance, physiological responses to varying tasks, and visual attention focus and patterns. Additional sensor data provided beyond this subset can further support research in IoT and home assistive robotics. For all experiments, we adopt a cross-subject evaluation method. Individual subjects are exclusively assigned to either the train or test set, thereby ensuring that no subject’s data is used in both training and testing. To maintain consistency across all experiments, the sets of subject IDs used for training and testing remain fixed. For reproducibility and further customization, our specific train/test subject split configuration is provided as a Python script in our GitHub repository  [^1] https://github.com/olivesgatech/DARai. For all classification tasks, we use class-wise top-1 accuracy. The class-wise accuracy metric is chosen because the hierarchical structure of human activities does not necessarily include all actions or procedures in every instance, resulting in a nonuniform distribution of lower-level classes. More details about the class distribution can be found in Appendix LABEL:Appendix:_Annotation.
We highlight the key observations from our data collection, sensor setup and each experiment in Table LABEL:tab:key_observations. A more comprehensive version of this table is provided in Appendix LABEL:Appendix:_observation-challenges.
Table 4: Key observations
Topic Key Findings
Visual Model Training Strategies Finetuned visual models consistently outperform those trained from scratch, across all class categories and all three levels of hierarchy.
Within kitchen and living room subsets, depth models trained from scratch outperform their RGB counterparts.
At levels L2 and L3 of the hierarchy, depth models continue to outperform RGB models trained from scratch, likely due to the stability of depth-based spatial representations across hierarchy levels.
Environment Differences Activity recognition accuracy within the living room is consistently higher than those within the kitchen.
Viewpoint and Placement Effect Visual modalities suffer a severe drop in accuracy when tested on the cross view.
Non-visual modalities are robust when the sensors capture symmetric interaction such as insole.
Wearable cross-body models show less performance degradation when compared to cross-view camera models.
Unimodal Sensor Strengths Gaze data are effective at separating fine-grained activities when other data modalities that rely on physical movements suffer from performance degradation.
Bio-monitoring data are good at identifying changes in physiology, such as increased heart rate in Playing video games or reduced activity in Sleeping. However, they lack the resolution to separate activities with similar effort, such as different cooking activities.
Hierarchical Granularity Changes in performance across granularity levels are not linear.
[…]
Insole pressure and Hand EMG data modalities do not exhibit a significant decline in performance from L1 to L2 as visual data does.
All modality combinations experience a decline in accuracy from L1 to L3, but the magnitude of degradation varies significantly by sensor modality combination.
Gaze data improves L3 (procedure) activity recognition fused with other modalities.
Biomonitoring signals do not offer fine-grained insights at Level 3, even when combined with insole, EMG, and IMU data modalities.
Temporal Localization In temporal activity localization, both the camera view and the length of the input video play a crucial role in accurately segmenting an untrimmed activity sample into lower-level segments.
Actions and procedures are not isolated events. They are interconnected components of larger sequences, with a strong dependence on the temporal context of those that come before and after.
Short- and  Long-Horizon Anticipation At L2 (Action level), a moderate observation period provides sufficient temporal context for immediate action anticipation, while excessive observation may introduce redundant or less relevant information.
Long-horizon anticipation at the L2 level, however, remains consistently lower and exhibits minimal improvement as the observation rate increases. This indicates that action-level predictions do not benefit significantly from extended temporal context, possibly due to the more independent nature of individual actions.
At the L3 (Procedure level), short-horizon anticipation improves as the observation rate increases, reaching optimal performance around a 70% observation rate.
Long-horizon anticipation at this level consistently improves with increasing observation rates, eventually surpassing short-horizon performance beyond a 70% observation rate. This trend suggests that fine-grained procedural activities exhibit stronger temporal dependencies, benefiting from longer observation windows that provide richer contextual information.
5 Visual Data Robustness and Real-World Limitations
Visual sensing modalities, such as RGB and depth, are effective for human activity recognition (Ma et al., 2019). However, such modalities suffer from robustness challenges in uncontrolled settings (Hara et al., 2021; Ponbagavathi et al., 2024). In particular, existing benchmarks that rely on vision modalities are generally limited to coarse-grained actions and do not address real-life scenarios. In contrast, DARai, with diverse modalities and multiple levels of hierarchical labels, was designed and curated to provide a benchmark for real-life deployment. This section presents a benchmark for evaluating the robustness of visual activity recognition models across varying viewpoints and environments.
5.1 Visual Data Benchmarking
Given the rise of large-scale pretrained models, we examined whether pretraining compensates for the diversity of daily activities. We compared models trained from scratch on DARai with models first pretrained on Kinetics 400 and then fine-tuned on DARai. For this benchmark, we evaluated R3D (Tran et al., 2018), MViT-small (Li et al., 2022), and Swin-tiny (Liu et al., 2022). Table LABEL:tab:visual-only-l1 in Appendix LABEL:Appendix:_ablation_tables compares their top-1 accuracy on combined and individual environment classes, while Table 5 reports performance by camera view at hierarchy levels L2 and L3. Although accuracy is high at the activity level (L1), all models show a notable drop at finer-grained levels (L2, L3).
Class Category All
Data Camera View Level Model Finetuned Trained from scratch
RGB $1$ L2 ResNet 0.46 $\pm$ 0.006 0.31 $\pm$ 0.01
$2$ 0.39 $\pm$ 0.01 0.25 $\pm$ 0.02
$1$ Mvit-s 0.43 $\pm$ 0.032 0.07 $\pm$ 0.004
$2$ 0.37 $\pm$ 0.015 0.06 $\pm$ 0.015
$1$ Swin-t 0.40 $\pm$ 0.007 0.16 $\pm$ 0.006
$2$ 0.31 $\pm$ 0.7 0.10 $\pm$ 0.027
$1$ L3 ResNet 0.52 $\pm$ 0.006 0.38 $\pm$ 0.01
$2$ 0.45 $\pm$ 0.007 0.28 $\pm$ 0.01
$1$ Mvit-s 0.53 $\pm$ 0.01 0.14 $\pm$ 0.07
$2$ 0.44 $\pm$ 0.02 0.05 $\pm$ 0.003
$1$ Swin-t 0.47 $\pm$ 0.02 0.25 $\pm$ 0.02
$2$ 0.39 $\pm$ 0.01 0.14 $\pm$ 0.016
Depth $1$ L2 ResNet - 0.60 $\pm$ 0.01
$2$ 0.53 $\pm$ 0.01
$1$ Mvit-s 0.32 $\pm$ 0.02
$2$ 0.18 $\pm$ 0.06
$1$ Swin-t 0.41 $\pm$ 0.04
$2$ 0.40 $\pm$ 0.008
$1$ L3 ResNet - 0.42 $\pm$ 0.01
$2$ 0.38 $\pm$ 0.01
$1$ Mvit-s 0.166 $\pm$ 0.03
$2$ 0.07 $\pm$ 0.04
$1$ Swin-t 0.21 $\pm$ 0.03
$2$ 0.24 $\pm$ 0.02
Table 5: Top-1 accuracy results for visual data experiments using three model architectures on Action and Procedure level of DARai hierarchy. Since activities across subsets may share similar actions or procedures, they cannot be separated by activity class subset. Results are reported for all available actions and procedures classes. Fine-tuned experiments are performed using models pretrained on Kinetics 400 dataset.
Analyzing these results presented in Table LABEL:tab:visual-only-l1 and Table 5 reveals a number of key observations:
Finetuned vs. Trained from scratch
Pretrained models consistently outperform those trained from scratch at all hierarchy levels. This advantage is most evident when training data are limited, where fine-tuned models leverage diverse prior features for higher accuracy.
Environment-Specific Activities
Recognition accuracy in the living room is consistently higher than in the kitchen. Even after fine-tuning on kitchen data, pretrained models still maintain a notable performance gap, suggesting stronger alignment with living-room-style activities.
Fine-Grained Levels (L2 and L3)
At L2 and L3, trained from scratch depth models surpasses RGB counterparts, likely due to stable spatial representations. However, once using pretrained weights, RGB models achieve higher performance.
5.2 Robustness to Real-World Conditions
5.2.1 Visual Cross-View Evaluation
DARai provides real-world conditions to test model robustness. For example, Figure 6 shows cross-view evaluations on three models, namely ResNet, MViT-s, and Swin-t, trained on a camera view and tested on same-view versus cross-view using RGB and depth data, respectively. As expected, these results show that models suffer a severe drop in accuracy when tested on the cross view. Objects and subjects that are fully within the viewpoint of a camera may not necessarily be viewable by the other camera. Thus, a significant drop in performance, e.g., ResNet performance drops from  89% to  15% and same model trained on the second view accuracy drops from  72% to 5%, is expected. Interestingly, depth data shows a similar trend although depth provides view-invariant information such as distances. This can be explained by the occlusion across views. Overall, vision-based recognition is tightly coupled to the camera perspective, a serious drawback for daily life applications within a home environment.
Figure 6: Comparison of same-view and cross-view inference performance for ResNet, MViT-s, and Swin-t using visual data. The results show top-1 accuracy for each model when tested on the same camera view and cross views.
5.2.2 Wearable Cross-Body Evaluation
While cross views are typical for RGB and depth modalities, wearable sensors have a cross-body tradeoffs. In this case, data from a sensor worn on one side of the body, e.g., a hand, an arm, or a foot, is used for training while testing is performed on data from the same type of sensor on the other side of the body. This setup creates a “cross-body” domain shift. In such settings, the challenges are different from those in vision-based cross-view systems. Certain body activities are mostly symmetric across the left and right parts of the body. Those activities are generally related to movement patterns and balancing body weights and pose, represented by foot-pressure insoles as depicted in Figure 7 for the Insole sensors. In contrast, hands interactions are generally asymmetric and one hand is usually dominant. As a result, a model trained on right-wrist or right-hand signals may struggle when tested on data from the left side, which can have a weaker or distinct signal, leading to higher error rates. This is clearly depicted in Figure 7 for the IMU and EMG sensors.
Figure 7: Comparison of same-view and cross-view inference performance using IMU, forearm muscle EMG, and insole pressure wearable data. The results show top-1 accuracy for each sensor when tested on the same side of the body and the opposite side. Models were trained separately on left-side and right-side data for this experiment.
In general, based on DARai, wearable cross-body models show less performance degradation when compared to cross-view camera models. This observation leads to the question about the contribution of every modality to recognition at every hierarchy. There seems to be a collaboration across these modalities to create a representation space that outperforms the counterparts created by a single or a subset of these modalities. DARai provides the proper settings to conduct such investigation.
6 Multimodal and Hierarchical Human Activities Benchmark
In this section, we evaluate unimodal and multimodal approaches on DARai across the three hierarchical levels (L1, L2, L3). We employ a transformer-based architecture, introduced in Appendix LABEL:Appendix:_ablation_tables, to analyze each modality alone and in pairwise combinations. Figure 8 presents these experiments results, highlighting both individual performance differences and the advantages of fusing complementary data modalities.
(a) A transformer-based architecture was trained from scratch on each modality separately for this experiment, using Swin-t for visual data modalities and the modified transformer introduced in Appendix LABEL:Appendix:_ablation_tables for other data modalities.
(b) A transformer-based architecture was trained to fuse each modality pairs separately for this experiment, using Swin-t for visual data modalities and the modified transformer introduced in Appendix LABEL:Appendix:_ablation_tables for other data modalities.
Figure 8: Comparison of Top-1 accuracy results for individual data modalities and data modality pairs across 3 level of DARai hierarchy.
6.1 Unimodal Analysis and Comparison
Table 6 summarizes the effectiveness of different sensing modalities in recognizing human activities. Each modality excels in specific contexts but also has limitations. For example, gaze data is effective for activities involving sustained visual fixation, while EMG signals capture distinct muscle activation patterns. Biomonitoring data provide insights into physiological changes but lack fine-grained resolution for differentiating similar exertion levels. Given such complementary effectiveness, it is common to fuse these modalities to improves accuracy and robustness. A detailed study on fusion based on DARai is in Appendix LABEL:Appendix:_ablation_tables and Figure LABEL:fig:unimodal-cm.
Modality Best Suited For Example Activities
Gaze Sustained visual fixation Reading, Watching TV, Working on a computer
EMG Muscle activation patterns Making pancakes (whisking motion), Playing video games (controller grasping)
Bio-Monitoring Physiological variations Playing video games (increased heart rate), Sleeping (reduced activity)
Insole Planter pressure patterns Carrying objects (added weight changes foot pressure), Standing postures
IMU Dynamic body movements Exercising (dynamic motion), Working on a computer (static)
Table 6: Comparison of Different Modalities for Activity Recognition
DARai’s unique hierarchical and multimodal aspects provide a space to investigate modality contribution to every level of the hierarchy. Figure 8(a) shows the performance of each of the modalities. The best-performing unimodal models for non-visual data modalities are shown on this plot. The input length, window size, and other parameters are kept consistent across all levels of hierarchy, with samples drawn exclusively from one hierarchy level at a time (L1, L2, or L3). As shown in Figure 8(a), changes in performance across granularity levels are not linear. Furthermore, different modalities respond differently to decreasing abstraction, shorter sample sequences, and finer-grain details. Previous studies have shown that task granularity negatively impacts natural image models (Prabhushankar et al., 2020; Kokilepersaud et al., 2024). Our findings confirm that even the visual model with the highest accuracy (i.e., 56%-59%), experiences a significant performance drop when transitioning from L1 to L2 and L3. In contrast, insole pressure, forearm muscle EMG and biomonitoring modalities do not exhibit the same steep decline in performance, experiencing only 5%, 2%, and 3% drops from their L1 accuracy, respectively. Notably, the decline in visual model performance is smaller between L2 and L3, with both camera views performing better at L3 than at L2. This trend is expected, as 20% of the procedures (L3) are shared between higher-level activities (L1) and actions (L2), indicating more overlap at these levels.
In contrast, biomonitor signals and hand muscle EMG signals demonstrate greater resilience to increased granularity between L1 and L2. Additionally, insole pressure outperforms other non-visual modalities at the action level (L2) with 32% accuracy. The alignment between plantar patterns and the structure of actions makes insole data effective for recognizing fine-grain actions at this level. However, from L2 to L3, none of these modalities maintain the same level of resilience. In naturalistic daily life activities, data are not trimmed around a single action, so wearable time-series often include extended inactive intervals, such as minimal hand movement or prolonged standing in one signal, while other modality signals remain more active. These intervals can overshadow distinct activity patterns within a single modality and lead to performance decline. At lower levels of the hierarchy, the increased granularity magnifies this effect, making it harder to capture pattern differences.
6.2 Multimodal Analysis and Comparison
We examine the performance of multimodal modeling across the three hierarchical levels of DARai. Specifically, we evaluate each modality combination detailed in Table 5, and summarize the findings in Figure 8(b).All modality combinations exhibit a performance decrease from L1 to L3, though the extent of this decline varies considerably between combinations.For instance, RGB+D performance decreases from 59% at L1 to 21% at L3, reflecting its difficulty in consistently capturing fine-grained distinctions. In contrast, fusing IMU, Insole, and gaze modalities results in a decrease from 57% at L1 to 41% at L3, suggesting that integrating motion, Insole pressure, and gaze data helps maintain the distinguishing capability of multimodal models. Combining all on-body sensors with gaze achieves 42% accuracy at L3, among the highest at this level. It outperforms more body-part-specific pairings, such as EMG with biomonitor data or Insole with biomonitor data, which yield significantly lower accuracy at L3 (38.9% and 7%, respectively). Some sensor combinations show a sharp accuracy drop from L1 to L3. For example, EMG-Insole accuracy decreases from 40% to 7%, and IMU-biomonitor fusion falls from 46.9% to 12.48%. In contrast, using all on-body sensors or fusing IMU, Insole, and gaze leads to higher L3 accuracy (23% and 41.2%, respectively) while biomonitors do not provide granular insights.
7 Temporal Dependencies Benchmark
Understanding temporal dependencies is crucial for modeling human activities, which unfold hierarchically over time. Unlike existing datasets with predefined step boundaries, DARai presents a continuous, unstructured setting where activities transition fluidly across different granularities, posing challenges in recognizing hierarchical structure and temporal dependencies. To assess these dependencies, we focus on temporal activity localization and activity anticipation. Performance is evaluated using Mean over Classes (MoC), which measures per-class performance and averages across all classes, following the long-term action anticipation protocol (Gong et al., 2022; Abu Farha et al., 2021; Sener et al., 2020; Ke et al., 2019). For consistency, results are averaged over three different seeds. For the activity localization result, refer to LABEL:sec:app-localization.
Figure 9: Feeding different length of sample sequence for action anticipation task. (a) shows short-term action anticipation task, and (b) shows long-term action anticipation task.
7.1 Short- and Long-Horizon Activity Anticipation
Action anticipation predicts future actions from partially observed sequences. DARai presents extended, fine-grained actions and procedures from a third-person view, enabling two tasks: short-horizon, which focuses on immediate upcoming actions, and long-horizon, which targets more distant actions. Figure 9 illustrates the setup, and we vary the observation ratio $\alpha$ from 0.1 to 0.9.
Figure 10: Action anticipation accuracy across different observation rates based on partially observed L1-level sequences at L2 (Action level) and L3 (Procedure level). The left graph represents L2-level action anticipation for both anticipating over the next 8 seconds (short-term action anticipation) and the latest 8 seconds of each sequence (long-term action anticipation). The right graph represents L3-level action anticipation for both short-term and long-term action anticipation. The accuracy represents Mean over Classes (MoC) accuracy.
Figure 10 shows that short-horizon anticipation improves steadily with more observation—up to about 50%—but gains little beyond that. This indicates that a moderate window captures enough context, and extra observations can introduce redundancy. In contrast, long-horizon anticipation consistently benefits from increasing the observation ratio, reflecting the strong temporal dependencies of extended procedural activities in DARai.
7.2 Hierarchical Activity Anticipation
Finer-grain anticipation tasks add additional complexity, as the model must understand both the hierarchical structure and the sequence of events or actions from higher-level activities. This type of anticipation is far more challenging than traditional feature prediction based on observed features. To address such challenges, DARai is essential for testing and benchmarking models, since it provides multiple levels of hierarchy.
Figure 10 shows the disparity in anticipation performance across different levels of temporal granularity. At the action level (L2), short-term action anticipation accuracy fluctuates between 42% and 50%, while long-term anticipation remains lower, ranging from 20% to 25% across different observation rates. This suggests that action anticipation primarily relies on local temporal cues rather than extensive historical context. Conversely, at the procedure level (L3), long-term anticipation accuracy at L3 improves from 36% to 46%, while short-term varies more significantly, ranging from 26% to 34% as observation rate changes. These findings emphasize the importance of considering the hierarchical nature of activities when designing anticipation models, as the required observation window and predictive capability vary depending on the level of granularity. Thus, optimizing anticipation strategies should be guided by the underlying temporal structure of the target activities.
8 Conclusion
We introduce the DARai dataset, a hierarchical, multimodal resource capturing everyday human activities in continuous, unscripted recordings. We highlight the following key observations. Detailed observations are summarized in Table LABEL:tab:key_observations.
Daily activities are not action-filled. Unlike datasets centered around actions, DARai has long periods of inactivities that includes thinking and searching, across all annotation hierarchies.
Unscripted activities promote counterfactual data collection. Providing activity categories and not providing scripts to participants significantly alters the temporal relationships between action and procedure segments.
Annotating continuous recordings of real-world human activities is challenging. Transitions between actions lack clear boundaries, complicating continuous human activity annotation. DARai provides hierarchical annotations that capture this fluidity and reflect inherent variability, enabling hierarchical modeling of activities in natural settings.
Hierarchical detail matters. Coarse activities are easier to classify than fine-grain actions and procedures.
Models can not handle placement shifts robustly. Models trained on a specific camera angle or body side drop in accuracy when tested on a different view or limb, underscoring the need for flexible, adaptive methods.
Sensor fusion enhances performance. No single sensor excels in all scenarios. Fusing wearables and gaze data improves performance even at finer levels of the hierarchy.
Temporal context impacts localization and anticipation. Actions have greater temporal diversity and complexity than procedures. Consequently, longer observation windows help localize actions more effectively, while predicting long-horizon procedures is generally easier than actions.
Limitations
The large sensor setup restricts flexible data collection and expansion, leading to the adoption of a hybrid approach. Future efforts could broaden environment coverage and explore fully wireless sensor setups. Additionally, our custom portable sensor system cost around \$$\$22,000$, making data-collection reproducibility a logistical issue.
Broader Impact
By showing how non-visual modality fusion improves recognition, DARai encourages privacy-focused approaches, such as minimal or no camera use. It also emphasizes the need for robust models that handle changes in placement and viewpoint. The dataset is released with user consent and promotes open-source AI and ML research.
9 Acknowledgments and Disclosure of Funding
This work is partially funded through the support from the Franklin Foundation via the John and Marilu McCarty Chair professorship and a gift from Lab126 at Amazon.